- Find the probability P(3 or 5) ,P(multiple of 2) ,P(number greater than 1) when a fair die is rolled. Thanks
Answer #1

If a fair die is rolled then number of possible outcomes are

n(s) = { 1, 2, 3 ,4 ,5 , 6 } = 6

1) probability of 3 or 5 is P(3 or 5 ) = P(3) + P(5)

P(3) = 1/6 and P(5) = 1/6

P(3 or 5 ) = (1/6)+(1/6) = 1/3

2) number outcomes of multiple of 2 = {2,4,6} = 3

P( multiple of 2 ) = number of outcomes multiple of 2 / n(s)

P( multiple of 2 ) = 3/6 = 1/2

3) P(greater than 1) = no. of outcomes greater than 1 / n( s)

P( greater than 1 ) = 5/6